# Two-Phase Channel Estimation for RIS-Assisted THz Systems with Beam Split

### v2

@@ V Numerical Results and Discussion
In this section, we present numerical results to validate the effectiveness of the proposed scheme. Without loss of generality, we consider a wideband RIS-assisted THz scenario with central carrier frequency $f_{c}=100$ GHz, bandwidth $B=15$ GHz and $M=128$ SCs. A BS employing a ULA with $N_{\mathrm{T}}$ = 16 antennas transmits signals to $K=8$ single antenna UEs, with the assistance of a RIS equipped with $N_{\mathrm{R}}=64$ passive reflecting elements. For both BS-RIS and RIS-UEs channels, there exit $L=J_{k}=3,\forall k$ propagation paths, and the AoAs and AoDs are randomly generated from the discretized grid within the range $\left[-\pi/2,\pi/2\right]$. The quantization level of overcomplete dictionaries is set as $Q_{\mathrm{T}}=32$ and $Q_{\mathrm{R}}=128$.
@@
We use the normalized mean square error (NMSE) as a metric for CSI estimation performance, which is given by
@@
$\mathrm{P}_{\mathrm{c}}=\mathrm{Pr}\left\{\left|\hat{\theta}-\theta\right|<\varepsilon_{\mathrm{th}}\right\},$ (50)
where $\varepsilon_{\mathrm{th}}\!=\!\delta/2$ is the threshold which is set as half the sample interval. For comparison, we also consider the conventional OMP scheme [30], BSA-OMP scheme [17], CBS-GAMP scheme [23], and oracle-LS scheme [31] as benchmark schemes in the simulation. It is noted that the angular information is assumed perfectly known in oracle-LS scheme, which can be regarded as the performance upper bound of the proposed scheme.
@@
Fig. 6 depicts the correct probability performance of the DoDs at BS and coupled DoAs/DoDs at RIS versus SNR with $T=50$ and $P=30$. It is shown that both the correct probabilities approach one as SNR increases. This indicates that the proposed scheme can ensure a certain level of reliability even under low SNR conditions and has the ability to obtain accurate physical directions with sufficient channel conditions. Additionally, the correct probability of the coupled directions at RIS is consistently lower than that of the DoDs at BS across all SNR values. This is because the estimation of the coupled directions at RIS depends on the accuracy of the DoDs estimation at BS. Angular estimation errors at BS would propagate to RIS, leading to performance gap between the two probability curves.
@@
Fig. 7 illustrates the NMSE performance versus SNR with $T=50$ and $P=30$. Overall, it can be observed that the proposed scheme outperforms conventional OMP and BSA-OMP algorithm across all SNR regions. This is because the proposed scheme has the ability to exploit unknown channel information, and the proposed dictionaries can effectively cover the beam space under beam split effect. Additionally, the proposed scheme is slightly inferior to that of CBS-GAMP algorithm when SNR is less than $10$ dB. This is because estimation errors generated by partial SCs propagate more significantly to the remaining SCs under poor channel conditions. However, the proposed scheme outperforms CBS-GAMP algorithm in high SNR regions, i.e., SNR is larger than $10$ dB. This is because angle decoupling provides more accurate angular estimation information. This suggests that there exists a performance-complexity trade-off between these two algorithms. Moreover, it is obvious that the proposed scheme can approach the NMSE performance of oracle-LS scheme, which verifies the effectiveness of the proposed scheme.
@@
Fig. 8 shows the NMSE performance versus the number of subframes $T$, with SNR $=20$ dB and $P=15$. As expected, the NMSE performance of all the algorithms improves with the increase in the number of subframes $T$. It can be observed that the proposed scheme maintains good performance even with a short length of subframes, e.g., $T=10$, demonstrating the robustness of the proposed algorithm. Moreover, compared to the CBS-GAMP method, the proposed algorithm achieves similar performance with approximately a $50\%$ subframes reduction. This indicates that the proposed scheme can efficiently reduce the pilot overhead for channel estimation.

### v3

@@ VII Numerical Results and Discussion
In this section, we present numerical results compared with both mmWave and THz benchmarks to validate the effectiveness of the proposed scheme. Without loss of generality, we consider wideband RIS-assisted communication scenarios with $M=128$ SCs. A BS employing a ULA with $N_{\mathrm{T}}$ = 16 antennas transmits signals to $K=8$ single antenna UEs, with the assistance of a RIS equipped with $N_{\mathrm{R}}=64$ passive reflecting elements. Moreover, the hybrid beamformer vector $\mathbf{w}_{t,p}[m]$ is assumed to be randomly distributed [17].
VII-A Comparison with mmWave Solution
In this subsection, we present simulation results to illustrate the distinct characteristics and impacts of beam squint and beam split effects, and further demonstrate the unique advantages of the proposed framework in comparison with mmWave solutions.
VII-A1 Setup
For the mmWave system, the central carrier frequency is set to $f_{c}=28$ GHz with a bandwidth of $B=600$ MHz, following the configuration in [23]. For the considered THz system, the central carrier frequency is increased to $f_{c}=100$ GHz with a wider bandwidth of $B=15$ GHz. Furthermore, the channel characterization adopts the same parameters as those specified in [24], and the path gains are model as $\alpha_{l}\sim\mathcal{CN}(0,1)$ and the time delays are constrained by $\tau_{l}\leq\tau_{\max}$ with $\tau^{\mathrm{mmWave}}_{\max}=53$ ns and $\tau^{\mathrm{THz}}_{\max}=20$ ns for mmWave and THz systems, respectively.
Additionally, we assume that the DoDs at BS are available as prior knowledge in this subsection, since the considered benchmark focused only on estimating the angles at RIS. Accordingly, the number of propagation paths is set to $L=1$ and $J_{k}=2,\forall k$, respectively. Furthermore, the DoAs and DoDs are randomly generated within the range $\left[-\pi/2,\pi/2\right]$. To further highlight the effect of beam squint and beam split, we consider the DoDs at RIS to be under large-angle reflection conditions, which significantly deviate from the normal incidence.
VII-A2 Benchmark
We adopt the NOMP based mmWave solution proposed in [23] as the benchmrk method. This approach is implemented for both mmWave and THz band, referred to as ‘NOMP-mmWave’ and ‘NOMP-THz’, respectively.
VII-A3 Metric
The root mean square error (RMSE) is utilized to evaluate the accuracy of angle estimation, defined as
$\mathrm{RMSE}=\sqrt{\mathbb{E}\left\{\left|\theta-\hat{\theta}\right|\right\}},$ (47)
where $\theta$ and $\hat{\theta}$ represent the true angle and the estimated angle, respectively.
VII-A4 Performance Analysis
For a fair comparison, identical pilot resources are allocated to both the NOMP-based solution and the proposed approach for estimating cascaded angles at RIS. Let $M_{\mathrm{sub}}$ denote the number of SCs used for angle estimation. The resource parameters of NOMP algorithm are set to $T=50$ and $M_{\mathrm{sub}}=12$, with the hybrid beamforming functionality disabled, consistent with the original implementation settings in [23]. Accordingly, the proposed method utilizes $T=25$, $P=12$ and $M_{\mathrm{sub}}=2$ resource elements.
Fig. 5 shows the RMSE performance of the coupled DoAs/DoDs estimation at RIS. While the NOMP-based mmWave solution achieves satisfactory performance in conventional mmWave scenarios, it fails to adequately address the beam split effect in THz systems. This degradation is primarily attributed to the substantially larger SC spacing in THz band, which induces more severe beam misalignment compared to the relatively mild beam squint effect observed at mmWave frequencies. As a result, conventional mmWave solutions fail to mitigate this misalignment by merely accumulating information from multiple SCs. In contrast, the proposed approach integrates a specifically designed DS-MUSIC algorithm, which distinguishes the true angular components from the frequency-dependent false ones by leveraging the inter-SC misalignment within an extended angular range, thereby achieving significantly enhanced estimation accuracy. Moreover, increasing the angular dictionary size $Q_{\mathrm{R}}$ leads to additional performance gains by mitigating quantization errors in the angle domain.
Figure 5: RMSE performance of coupled DoAs/DoDs at RIS against SNR.
VII-B Comparison with THz Solutions
In this subsection, we investigate the impact of different system parameters on performance, and provide a comprehensive comparison with state-of-the-art algorithms in THz band.
VII-B1 Setup
We consider the THz system as defined in Section VII-A. Without loss of generality, the DoAs and DoDs are randomly generated from the discretized grid within the range $\left[-\pi/2,\pi/2\right]$ in this subsection [18]  [^8] The quantization errors introduced by discrete angular grids assumption are small for large dictionary quantization level $Q_{\mathrm{T}}$ and $Q_{\mathrm{R}}$, which do not affect significantly the normalized array gain and can be neglected.. Consequently, the quantization level of overcomplete dictionaries can be reduced to $Q_{\mathrm{T}}=32$ and $Q_{\mathrm{R}}=128$ for simplicity, and the 1-st and 64-th SCs are selected for the full CSI estimation in Phase I Step 1. Furthermore, we consider a more practical multipath model where the location of both BS and RIS is unknown and there exists $L=J_{k}=3,\forall k$ propagation paths for both BS-RIS and RIS-UEs channels.
VII-B2 Benchmarks
For comparison, we also evaluate the following benchmark schemes in the simulation. Unless otherwise specified, they are configured identically to the proposed method to ensure a fair comparison.
  • Conventional-OMP [35]: Traditional dictionaries are adopted to span the beam space and the CSI is reconstructed by OMP algorithm at each SC.
  • BSA-OMP [17]: Beam-split-adapted dictionaries are employed to cover the beam space and the CSI is recover by OMP algorithm at each SC.
  • CBS-GAMP [26]: CBS dictionaries are employed to cover the beam space and the CSI is recover by GAMP algorithm at each SC.
  • Oracle-LS: The angular information is assumed perfectly known and the angle-excluded information is estimated by LS algorithm. This algorithm is regarded as the performance upper bound of the proposed scheme.
VII-B3 Metrics
@@
We use the normalized mean square error (NMSE) as a metric for full CSI estimation performance, which is given by
@@
$\mathrm{P}_{\mathrm{c}}=\mathbb{E}\left\{\mathrm{Pr}\left[\left|\theta-\hat{\theta}\right|<\varepsilon_{\mathrm{th}}\right]\right\},$ (49)
where $\varepsilon_{\mathrm{th}}\!=\!\delta/2$ denotes the threshold which is set as half the sampling interval.
VII-B4 Performance Analysis
@@
Fig. 6 depicts the correct probability performance of the DoDs at BS and coupled DoAs/DoDs at RIS versus SNR with $T=50$ and $P=30$. It is shown that both the correct probabilities approach one as SNR increases. This indicates that the proposed scheme can ensure a certain level of reliability even under low SNR conditions and has the ability to obtain accurate physical directions with sufficient channel conditions. Additionally, the correct probability of the coupled directions at RIS is consistently lower than that of the DoDs at BS across all SNR values. This is because the estimation of the coupled directions at RIS depends on the accuracy of the DoDs estimation at BS. Angular estimation errors at BS would propagate to RIS, leading to performance gap between the two probability curves. This discrepancy highlights the propagation of errors from Phase I to Phase II: the angular estimation errors at BS directly impact the estimation accuracy of the coupled directions at RIS, which in turn affects the CSI estimation for other SCs in subsequent steps. However, as the SNR increases, the influence of these propagated errors diminishes, which enables Phase II to achieve improved CSI recovery performance.
@@
Fig. 7 illustrates the NMSE performance versus SNR with $T=50$ and $P=30$. Overall, it can be observed that the proposed scheme outperforms conventional OMP and BSA-OMP algorithm across all SNR regions. This is because the proposed scheme has the ability to exploit unknown channel information, and the proposed dictionaries can effectively cover the beam space under beam split effect. Additionally, the proposed scheme is slightly inferior to that of CBS-GAMP algorithm when SNR is less than $10$ dB. This is because estimation errors generated by partial SCs propagate more significantly to the remaining SCs under poor channel conditions. However, the proposed scheme outperforms CBS-GAMP algorithm in high SNR regions, i.e., SNR is larger than $10$ dB. This is because angle decoupling provides more accurate angular estimation information. This suggests that there exists a performance-complexity trade-off between these two algorithms. Moreover, it is observed that the proposed scheme progressively approaches the NMSE performance of oracle-LS benchmark as SNR increases. Under high SNR conditions, the performance gap between the two schemes becomes negligible, further verifying the effectiveness of the proposed scheme.
@@
Fig. 8 shows the NMSE performance versus the number of subframes $T$, with SNR $=20$ dB and $P=15$. As expected, the NMSE performance of all the algorithms improves with the increase in the number of subframes $T$. It can be observed that the proposed scheme maintains good performance even with a short length of subframes, e.g., $T=20$, demonstrating the robustness of the proposed algorithm. Moreover, compared to the CBS-GAMP method, the proposed algorithm achieves similar performance with approximately a $50\%$ subframes reduction. This indicates that the proposed scheme can efficiently reduce the pilot overhead for channel estimation.
@@ VIII Conclusion
Appendix A Proof of Proposition 1
We first define the parameter sequence $\bar{\vartheta}_{q_{1}}=\delta\cdot q_{1}-(Q_{\mathrm{R}}+1)/{Q_{\mathrm{R}}},\forall q_{1}=1,\dots,Q_{\mathrm{R}}$ which corresponds one-to-one with $Q_{\mathrm{R}}$ ARVs in the dictionary $\mathbf{C}_{\mathrm{R},m}$. Similarly, we define the sequence $\delta_{q_{2}}=-\delta\cdot q_{2}+(Q_{\mathrm{R}}+1)/Q_{\mathrm{R}}$ to correspond to $Q_{R}$ ARVs in the conjugate dictionary $\mathbf{C}_{\mathrm{R},m}^{*}$. Thus, the column vectors of $\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}$ can be expressed as $\mathbf{a}_{\mathrm{R},m}(\bar{\vartheta}_{q_{1}}-\bar{\vartheta}_{q_{2}})$ for $q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$. The $j$-th column of the coupled dictionary $\tilde{\boldsymbol{\xi}}_{\mathrm{R},m,j}$ can be expressed as
where $j=\left(q_{2}-1\right)Q_{\mathrm{R}}+q_{1},\forall q_{1},q_{2}\in 1,\cdots,Q_{\mathrm{R}}$. Based on the fact that $q_{1}-q_{2}\in\left[1-Q_{\mathrm{R}},Q_{\mathrm{R}}-1\right]$, the number of unique columns of coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{\mathrm{B}}=2Q_{\mathrm{R}}-1$.
Figure 12: Coupled dictionary simplification diagram.
Then, we simplify the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ step by step from (a) to (d) as shown in Fig. 12. As depicted in step (a), we reshape the overcomplete dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ of length $Q_{\mathrm{R}}^{2}$ ARVs into a three-dimensional matrix where each column represents an ARV. Given that ARV can be uniquely determined by known angles in (50), we use the ordered sequence $\mathcal{L}_{Q}=\left(1,\dots,Q_{\mathrm{R}}\right)$ to index the corresponding ARVs. In step (b), we form the index matrix $\mathbf{Q}\in\mathbb{N}^{Q_{\mathrm{R}}\times Q_{\mathrm{R}}}$ with elements $[\mathbf{Q}]_{i,j}=(\mathcal{L}_{Q})_{i}-(\mathcal{L}_{Q})_{j}$ for $i,j=1,\dots,Q_{\mathrm{R}}$. It is noticed that due to the cyclic shift property of this dictionary, the dictionary base can be represented entirely by combining just the first and last rows. This means that we can reconstruct the original dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ with a sliding window of length $Q_{\mathrm{R}}$. Therefore, the simplified CBS dictionary $\mathbf{\Xi}_{\mathrm{R},m}$ can be given as in (14). Moreover, substituting $q_{2}=Q_{\mathrm{R}}$ into (50), the $q$-th column where $q=1,\dots Q_{\mathrm{B}}$ of $\mathbf{\Xi}_{\mathrm{R},m}$ can be formulated as